keq= 15 at 500 K. Predict how the reaction will proceed with following information?

2NOBr(g) (equilibrium) 2NO (g) + Br2(g)

A flask is charged with 4 atm of NO, 2 atm of Br2 and 2 atm of NOBr.
Answer #1

The Kp is calculated:

Kp = Kc * (R * T) ^ Δn = 15 * (0.082 * 500) ^ (3-2) = 615

The Qp is calculated:

Qp = P NO ^ 2 * P Br2 / P NOBr ^ 2 = 4 ^ 2 * 2/2 ^ 2 = 8

Since Qp <Kp, the reaction shifts to the right, forming products to reach equilibrium.

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.